OpenStudy (cailindevereux):

Use a graphing device to find all solutions of the equation, rounded to two decimal places. (Enter your answers as a comma-separated list.) 2^(−x) = x − 4

OpenStudy (sooobored):

uh, so use a graphing calculator and the build in zero function to find the roots thats what i think your question is asking for

OpenStudy (cailindevereux):

im not sure thats all the question says and 0,4 is incorrect

OpenStudy (sooobored):

was the original problem y=x^2 +x-4 ? cause you need to graph that not whatever contraption you have written down

OpenStudy (cailindevereux):

its exactly what i have written down

OpenStudy (cailindevereux):

and its negative 2 not positive

OpenStudy (cailindevereux):

*-x but i just copy and pasted the problem

OpenStudy (sooobored):

wait, so the problem looks like \[2^{-x}=x-4\]

OpenStudy (cailindevereux):

yes

OpenStudy (sooobored):

oh ok, just move the 2^-x over to the right side so youll get 0= -2^(-x)+x-4 and graph y=-2^(-x)+x-4 since finding the root means setting y=0 and solving for the solutions of x

Directrix (directrix):

I'm going to draw the graph and see what I get.

Directrix (directrix):

The instruction to find all solutions is what we need to think upon. I see the one real root on the graph but am unsure about how to interpret the general solution, @cailindevereux